Assume a par value of $1,000. Caspian Sea plans to issue a 13.00 year, annual pay bond that has a coupon rate of 7.93%. If the yield to maturity for the bond is 8.26%, what will the price of the bond be?

Assume a par value of $1,000. Caspian Sea plans to issue a 26.00 year, annual pay bond that has a coupon rate of 10.00%. If the yield to maturity for the bond is 10.0%, what will the price of the bond be?

Answer format: Currency: Round to: 2 decimal places.

The market price of a semi-annual pay bond is $953.09. It has 20.00 years to maturity and a coupon rate of 5.00%. Par value is $1,000. What is the yield to maturity?

Answer format: Percentage Round to: 4 decimal places (Example: 9.2434%, % sign required. Will accept decimal format rounded to 6 decimal places (ex: 0.092434))
Answer #1

1)

Coupon = 0.0793 * 1000 = 79.3

Price = Coupon * [1 - 1 /(1 + r)n] / r + FV/ (1 + r)n

Price = 79.3 * [1 - 1 /(1 + 0.0826)13] / 0.0826 + 1000/ (1 + 0.0826)13

Price = 79.3 * [1 - 0.356382]/ 0.0826 + 356.381954

Price = 79.3 * 7.791985 + 356.381954

Price = $974.29

You ca also find this using a financial calculator:

FV 1000

PMT 79.3

N 13

I/Y 8.26

CPT PV

2)

Price of bond = $1,000.00

When coupon rate is equal to yield to maturity, price will always be equal to par. Therefore, price will be $1000.

3)

Coupon = (0.05 * 1000) / 2 = 25

Number of periods = 20 * 2 = 40

Yield to maturity = 5.3860%

Keys to use in a financial calculator:

2nd I/Y 2

FV 1000

PV -953.09

PMT 25

N 40

CPT I/Y